The Petit Chef Co. has a 6% coupon bonds with 5 years left to maturity. The bonds make semi-annual payments and has a par value of $2,000. If the bonds sell for $2,200, what is the yield-to-maturity? Show all work
Answer #1

Face value of bond = $2000

coupon rate = 6% paid semiannual

So coupon = 3% of 2000 = $60 semiannual

total period = 2*5 = 10

Price = $2200

using financial calculator with values

FV = 2000

PV = -2200

PMT = 60

N = 10

compute for I/Y we get, I/Y = 1.893

So YTM = 2*1.893 = 3.786%